OpenStudy (sabrina1):

find the standard form of a circle with a diameter that has endpoints (-4,8) amd (9,-6)

hero (hero):

Find the midpoint of the diameter

OpenStudy (anonymous):

How do I do that?

OpenStudy (sabrina1):

thanks:)

OpenStudy (anonymous):

the midpoint of the diameter is the center\[x _{m}=.5(9-4)=5/2\]\[y _{m}=.5(8-6)=1\]the center of the circle is (5/2, 1)

OpenStudy (anonymous):

the circle follows the form\[(x-h)^2+(y-k)^2=r^2\]and we just found h and k so\[(x-5/2)^2+(y-1)^2=r^2\]now we just need to find the radius, or easier is the radius squared

OpenStudy (sabrina1):

yes radius is squared

OpenStudy (sabrina1):

should we just put in the diameter coordinates in equation of the circle and then solve to find radius?

OpenStudy (anonymous):

the radius is the distance from the center to either of the given points; i choose to use (-4,8); formula is \[r^2=(5/2-(-4))^2+(1-8)^2=\frac{169}{4}+49=\frac{365}{4}\]

OpenStudy (anonymous):

putting it all together\[(x-5/2)^2+(y-1)^2=\frac{365}{4}\]assuming i didn't make any arithmetic errors!